# Cohomogeneity one $\mathsf{RCD}$-spaces

### v2

@@ 5. Construction of cohomogeneity one $\mathsf{RCD}$-spaces
$=$ $-\frac{A\epsilon^{2}}{\sqrt{((Ax-1)^{2}+\epsilon^{2})^{3}}}\leq 0$
@@
Let $(F,\mathfrak{m}_{F})$ be a weighted Finsler manifold which satisfies the $\mathsf{CD}((N-1)K_{F},N)$-condition for $K_{F}\geq 0$ such that, if $N=1$ then $\operatorname{diam}(F)\leq\pi/\sqrt{K_{F}}$, and $\sqrt{K_{F}}\geq(A/2)(1+(1/\sqrt{1+\epsilon^{2}}))$. Then the $N$-warped product $[0,\infty)\times^{N}_{f_{A,\epsilon}}F$ is an $\mathsf{CD}^{\ast}(NK,N+1)$-space for $K=0$. Moreover, if $(F,\mathfrak{m}_{F})$ is infitesimally Hilbertian, then $[0,\infty)\times^{N}_{f_{A,\epsilon}}F$ is an $\mathsf{RCD}(NK,N+1)$-space for $K=0$.
@@
The interval $[0,\infty)$ with the standard Euclidean metric is an Alexandrov space of $\operatorname{curv}\geq 0$. By Lemma 5.14 a we have that $f_{A,\epsilon}$ is $\mathcal{F}K$-concave for $K=0$. Moreover we have $|f_{A,\epsilon}^{\prime}|\leq\sqrt{K_{F}}$. Thus the hypothesis of Theorem 5.13 are satisfied yielding the first claim. Moreover, when $(F,\mathfrak{m}_{F})$ is infinitesimally Hilbertian, we obtain the second conclusion from Proposition 5.12. ∎
With this result, we are able to show that spaces in Theorem A (b) admit an $\mathsf{RCD}(0,N)$-structure on which $G$ acts by measure preserving isometrics of cohomogeneity one.
@@
Let $(X,d,\mathfrak{m})$ be an $\mathsf{RCD}(K,N)$-space, and let $G$ be a compact Lie group acting on $X$ by measure preserving isometries, such that the orbit space $X/G$ is homeomorphic to $[0,\infty)$. Let $G_{0}$ be the unique singular isotropy group, and $H$ the principal isotropy group. Under the assumption that $G_{0}/H$ is simply-connected, there exists a new metric $\tilde{d}$ and measure $\tilde{\mathfrak{m}}$ on $X$ making into an $\mathsf{RCD}(0,N^{\prime})$-space, where $N^{\prime}=\dim(G_{0}/H)+1+\dim(G)$. Moreover, $G$ acts on $(X,\tilde{d},\tilde{\mathfrak{m}})$ by measure preserving isometries with cohomogeneity one.
@@
We consider on $G$ a bi-invariant Riemannian metric. By Lemma 4.8 and Theorem D, we have that cone fibers which are homeomorphic to the space $F=G_{0}/H$. As proven in [6], the restriction of the $G$-bi-inviariant Riemannian metric to $G_{0}$ induces a $G_{0}$ Riemannian metric of positive Ricci curvature on $F$. Thus the space $(F,d_{0},\mathfrak{m}_{0})$ with the induced metric and measure is an $\mathsf{RCD}(K_{0},N_{0})$-space with $K_{0}>0$, and $N_{0}=\dim(F)$. Thus there exists $K_{F}>0$ such that $K_{0}=(N_{0}-1)K_{F}$.
Observe that in the case when $N_{0}=1$, then by taking $\lambda\geq\operatorname{diam}(F)/(\pi/\sqrt{K_{F}})$, we have that by rescaling $F$ by $\lambda$, the rescaled space $(F,\lambda d_{0},(\lambda)^{N_{0}}\mathfrak{m}_{0})$ is an $\mathsf{RCD}((N_{0}-1)K_{F}/\lambda,N_{0})$-space which satisfies $\operatorname{diam}(F,\lambda d_{0})\leq\pi\lambda^{2}/\sqrt{K_{F}}$. Thus without loss of generality, we may assume that when $N_{0}=1$, we have $\operatorname{diam}(F)\leq\pi/\sqrt{K_{F}}$.
We consider now $A>0$ and $\varepsilon>0$ appropriately such that $\sqrt{K_{F}}\geq(A/2)(1+(1+\sqrt{1+\epsilon^{2}}))$. Then by Theorem 5.15 we have that the $N_{0}$-warped product $[0,\infty)\times_{f_{A,\epsilon}}F$ is an $\mathsf{RCD}(0,N_{0}+1)$-space. Thus the product metric measure space $G\times([0,\infty)\times_{f_{A,\epsilon}}F)$ is an $\mathsf{RCD}(0,N_{0}+1+D)$-space, where $D=\dim(G)$, and $G_{0}$ acts by measure preserving isometries via the diagonal action $\Delta\colon G\times([0,\infty)\times_{f_{A,\epsilon}}F)$ given by
@@
Thus by [27, Theorem 1.1] the quotient space $(G\times([0,\infty)\times_{f_{A,\epsilon}}F))/\Delta$ is an $\mathsf{RCD}(0,N_{0}+1+D)$ space. But by Theorem A this space is homeomorphic to $X$. ∎
We now consider $N$-warped products of the form $[0,2/A]\times_{f_{A}}^{N}F$:
@@
Let $(F,d_{F},\mathfrak{m}_{F})$ be a weighted Finsler manifold which satisfies the $\mathsf{CD}((N-1)K_{F},N)$-condition for $K_{F}\geq 0$ such that, if $N=1$ then $\operatorname{diam}(F)\leq\pi/\sqrt{K_{F}}$. Assume that $K_{F}\geq A^{2}$. Then the $N$-warped product $[0,2/A]\times_{f_{A}}^{N}F$ is a $\mathsf{CD}^{\ast}(0,N+1)$-space, and an $\mathsf{RCD}(0,N+1)$-space if $(F,d_{F},\mathfrak{m}_{F})$ satisfies the $\mathsf{RCD}((N-1)K_{F},N)$-condition.
@@
By the previous theorem, the $N$-warped products $[0,\infty)\times_{f_{A,\epsilon}}^{N}F$ are $\mathsf{CD}^{\ast}(0,N+1)$-spaces (respectively $\mathsf{RCD}(0,N+1)$-spaces if $F$ is infinitesimally Hilbertian).
We now consider the open subset $[0,2/A)\times F/(0,p)\sim(0,q)$. This is the open ball of radius $2/A$ centered at the vertex $\ast$ of the cones, with respect to the metrics $d_{A}$ or $d_{\epsilon}$. Moreover as we showed, it is totally geodesic, and by construction we have that $\mathfrak{m}_{\epsilon}(\partial[0,2/A)\times F/(0,p)\sim(0,q))=0=\mathfrak{m}_{A}(\partial[0,2/A)\times F/(0,p)\sim(0,q))$. Thus we conclude that $[0,2/A]\times_{f_{A,\epsilon}}F$ are $\mathsf{CD}^{\ast}(0,N+1)$-spaces (respectively, $\mathsf{RCD}(0,N+1)$-spaces when $F$ is infinitesimally Hilbertian).
Since the $N$-warped product $[0,2/A]\times_{f_{A}}^{N}F$ is the pointed measured Gromov-Hausdorff limit of the spaces $[0,2/A]\times_{f_{A,\epsilon}}^{N}F$ as $\epsilon\to 0$, and the $\mathsf{CD}^{\ast}$, $\mathsf{RCD}$-conditions are stable under this type of convergence, we obtain the desired result. ∎
@@
Theorem 5.17 ((9) in [58]).
If $(X,d,\mathfrak{m})$ is $\mathsf{CD}^{\ast}(K,N)$, then for $\lambda>0$ the space $(X,\lambda d,\mathfrak{m})$ is $\mathsf{CD}^{\ast}(\lambda^{-2}K,N)$.
We consider an $n$-dimensional closed Riemannian manifold $F$ such that for $0<\kappa_{F}$ we have $\kappa_{F}\leq\mathrm{Ric}(F)$, where $\mathrm{Ric}(F)=\min\{\mathrm{Ric}(x)\mid x\in TM\>\|x\|=1\}$. Thus we have that $(F,d_{F},\mathfrak{m}_{F})$ is a $\mathsf{CD}^{\ast}(\kappa_{F},n)$-space. By taking $\lambda=\sqrt{\kappa_{F}/4(n-1)}$ we have that $(F,\lambda d_{F},\mathfrak{m}_{F})$ is an $\mathsf{CD}^{\ast}((n-1)4,n)$-space. Given that $F$ is a Riemannian manifold, we obtain that $(F,\lambda d_{F},\mathfrak{m}_{F})$ is an $\mathsf{RCD}(0,n)$-space, and also $(F,\mathfrak{m}_{F})$ is a weighted Finsler manifold. Thus we have that $|f_{2,\epsilon}^{\prime}(0)|\leq 2$, and by Theorem 5.15 that $[0,\infty)\times_{f_{2,\epsilon}}(F,\lambda d_{F},\mathfrak{m}_{F})$ is an $\mathsf{RCD}(0,n+1)$-space. Since the balls of radius $1$ are geodesic, then by taking the limit as $\epsilon\to 0$ we obtain that that $[0,1]\times_{f_{2}}(F,\lambda d_{F},\mathfrak{m}_{F})$ is an $\mathsf{RCD}(0,n+1)$-space.
@@
We now consider gluings along the boundary of (certain quotients of) the $N$-warped products we have considered so far using the functions $f_{A}\colon[0,2/A]\to[0,\infty)$. Consider the cone-like warped product $C=[0,2/A]\times F/(0,p)\sim(0,q)$ with the distance $d_{C}=d_{A}$ and measure $\mathfrak{m}_{C}=f^{n-1}_{A}d\mathfrak{m}_{F}\otimes d\mathcal{L}$, where $(F,d_{F},\mathfrak{m}_{F})$ is an (infinitesimally Hilbertian) weighted Finsler manifold satisfying the $\mathsf{CD}^{\ast}((N-1)K_{F},N)$-condition for some $K_{F}>0$. We consider $G$ equipped with a bi-invariant metric $d_{G}$ and the Haar measure $\mathfrak{m}_{G}$. We take $G\times C$ equipped with the product distance $d_{G}\oplus d_{C}$ and measure $\mathfrak{m}_{G}\otimes\mathfrak{m}_{C}$. Thus, taking $D:=\dim(G)$, by Theorem 5.16 the space $G\times C$ is an $\mathsf{RCD}(0,N+1+D)$-space.
@@
Consider $G$ an $D$-dimensional compact Lie group, $K\subset G$ a Lie subgroup, and $(F,d_{F},\mathfrak{m}_{F})$ an infinitesimally Hilbertian weitghted Finsler manifold, which satisfies the $\mathsf{RCD}((N-1)K_{F},N)$-condition on which $K$ acts transitively by measure preserving isometries for some $K_{F}>0$. Assume that $\operatorname{diam}(F)\leq\pi/\sqrt{K_{F}}$, if $N=1$. Consider $A>0$ such that $K_{F}>A^{2}$, and the function $f_{A}(x):=\min\{1,Ax\}$. Consider the warped product $C:=[0,2/A]\times_{f_{A}}^{N}F$, and denote by $\Delta$ the diagonal action of $K$ on $G\times C$. Then the space $G\times C/\Delta$ is an $\mathsf{RCD}(0,N+1+D)$-space.
Moreover when we consider a homogeneous space $X=G/H$, equipped with $G$-invariant distance $d$ and measure $\mathfrak{m}$ making it an $\mathsf{RCD}$-space, the measure $\mathfrak{m}$ has to be (a multiple of) the Hausdorff measure as we now show (this has also been independently proved by [48, Theorem 2]).
@@
By [68, Proposition 5.14] we have that $X$ must be isometric to a Riemannian manifold. So, in particular we have that all the tangent spaces must be isometric to $\mathbb{R}^{n}$ for some natural number $n$. Then, by [12, Theorem 1.12] we have then that $\mathfrak{m}=\rho\mathcal{H}^{n}$, where $\mathcal{H}^{n}$ denotes the $n$-dimensional Hausdorff measure. We claim now that for all $g\in G$. $\rho\circ g=\rho$. Let $A\in\mathcal{B}(X)$, $g\in G$, and recall that $g_{\#}\mathcal{H}^{n}=\mathcal{H}^{n}$. As the action is by measure preserving isometries $\mathfrak{m}(A)=g_{\#}\mathfrak{m}(A)$. This implies
[…]
@@
Consider $G_{\pm}$ compact Lie groups, $(P_{\pm},d_{\pm}^{P},\mathfrak{m}_{\pm}^{P})$ be $N_{\pm}^{P}$-dimensional Riemannian manifolds with $\operatorname{Ric}(P_{\pm})\geq K^{P}_{\pm}$, and $(F_{\pm},d_{\pm}^{F},\mathfrak{m}_{\pm}^{F})$ be weighted Riemannian manifolds which are $\mathsf{RCD}(K_{\pm}^{F},N_{\pm}^{F})$ with $K_{\pm}^{F}>0$. Assume that $G_{\pm}$ acts effectively on $P_{\pm}$ and $F_{\pm}$ by measure preserving isometrties. Furthemore assume that the balanced product $P_{-}\times_{G_{-}}F_{-}$ is homeomorphic to $P_{+}\times_{G_{+}}F_{+}$. Then the glued space
@@
We choose $K_{\pm}>0$ so that $(F_{\pm},d^{F}_{\pm},\mathfrak{m}_{\pm}^{F})$ are $\mathsf{RCD}((N_{\pm}^{F}-1)K_{\pm},N^{F}_{\pm})$, for example by taking $K_{\pm}\leq K_{\pm}^{F}/(N^{F}_{\pm}-1)$. Observe that in the case when either $N^{F}_{+}=1$ or $N^{F}_{-}=1$, then by taking $\lambda_{\pm}\geq\operatorname{diam}(F_{\pm})/(\pi/\sqrt{K_{\pm}})$, we have that by rescaling $F_{\pm}$ by $\lambda_{\pm}$, the rescaled space $(F_{\pm},\lambda_{\pm}d^{F}_{\pm},(\lambda_{\pm})^{N^{F}_{\pm}}\mathfrak{m}^{F}_{\pm})$ is an $\mathsf{RCD}((N^{F}_{\pm}-1)K_{\pm}/\lambda_{\pm},N^{F}_{\pm})$-space which satisfies $\operatorname{diam}(F_{\pm},\lambda_{\pm}d^{F}_{\pm})\leq\pi\lambda_{\pm}^{2}/\sqrt{K_{\pm}}$. Thus without loss of generality, we may assume that if either $N^{F}_{+}=1$ or $N^{F}_{-}=1$, we have $\operatorname{diam}(F_{\pm})\leq\pi/\sqrt{K_{\pm}}$.
We now consider $A_{\pm}>0$ such that $K_{\pm}>A_{\pm}^{2}$. We now consider the following two $\mathsf{RCD}(0,N^{F}_{\pm}+1+N^{P}_{\pm})$-spaces:
@@
[…]

### v3

@@ 5. Construction of cohomogeneity one $\mathsf{RCD}$-spaces
$=$ $-\frac{A\epsilon^{2}}{\sqrt{((Ax-1)^{2}+\epsilon^{2})^{3}}}\leq 0.$
@@
Let $(F,\mathfrak{m}_{F})$ be a weighted Finsler manifold which satisfies the $\mathsf{CD}((N-1)K_{F},N)$-condition for $K_{F}>0$ such that, if $N=1$ then $\operatorname{diam}(F)\leq\pi/\sqrt{K_{F}}$, and $\sqrt{K_{F}}\geq(A/2)(1+(1/\sqrt{1+\epsilon^{2}}))$. Then the $N$-warped product $0,\infty)\times^{N}_{f_{A,\epsilon}}F$ is an $\mathsf{CD}^{\ast}(NK,N+1)$-space for $K=0$. Moreover, if $(F,\mathfrak{m}_{F})$ is infinitesimally Hilbertian, then $0,\infty)\times^{N}_{f_{A,\epsilon}}F$ is an $\mathsf{RCD}(NK,N+1)$-space for $K=0$.
@@
The interval $0,\infty)$ with the standard Euclidean metric is an Alexandrov space of $\operatorname{curv}\geq 0$. By ?THM? LABEL:L:_F_A_varepsilon_are_FK_concave a we have that $f_{A,\epsilon}$ is $\mathcal{F}K$-concave for $K=0$. Moreover we have $|f_{A,\epsilon}^{\prime}|\leq\sqrt{K_{F}}$. Thus the hypothesis of ?THM? LABEL:T:_Ketterer_warped_products are satisfied yielding the first claim. Moreover, when $(F,\mathfrak{m}_{F})$ is infinitesimally Hilbertian, we obtain the second conclusion from ?THM? LABEL:prop.infHilbertwarped. ∎
With this result, we are able to show that spaces in ?THM? LABEL:MC:_Homeomorphism_rigidity (b) admit an $\mathsf{RCD}(0,N)$-structure on which $G$ acts by measure preserving isometrics of cohomogeneity one.
@@
[…]
@@
We consider on $G$ a bi-invariant Riemannian metric. By ?THM? LABEL:L:_infinitesimal_action_cohomogeneity_1 and ?THM? LABEL:MT:_Geometry_of_the_slice, we have that cone fibers which are homeomorphic to the space $F=G_{0}/H$. As proven in [9], the restriction of the $G$-bi-invariant Riemannian metric to $G_{0}$ induces a $G_{0}$ Riemannian metric of positive Ricci curvature on $F$ when $\dim(F)\geq 2$. In the case when $N_{0}=1$, since $G_{0}$ and $H$ are compact ($G$ is compact and acts by isometries, thus the isotropy groups are compact), then $F$ is compact. Thus, there exists $K_{F}>0$ such that $\operatorname{diam}(F)\leq\pi/\sqrt{K_{F}}$, and $F$ with the metric induced by the restriction of the $G$-biinvariant metric to $G_{0}$ is an $\mathsf{RCD}(0,1)=\mathsf{RCD}((N_{0}-1)K_{F},N_{0})$-space, where $N_{0}:=\dim(F)$.
In case when $N_{0}\geq 2$, the space $(F,d_{0},\mathfrak{m}_{0})$ with the induced metric and measure is an $\mathsf{RCD}(K_{0},N_{0})$-space with $K_{0}>0$, and $N_{0}=\dim(F)$. Thus there exists $K_{F}>0$ such that $K_{0}=(N_{0}-1)K_{F}$.
We consider now $A>0$ and $\varepsilon>0$ appropriately such that $\sqrt{K_{F}}\geq(A/2)(1+(1+\sqrt{1+\epsilon^{2}}))$. Then by ?THM? LABEL:T:_cone_over_f_A_sigma_are_CD we have that the $N_{0}$-warped product $0,\infty)\times_{f_{A,\epsilon}}F$ is an $\mathsf{RCD}(0,N_{0}+1)$-space. Thus the product metric measure space $G\times([0,\infty)\times_{f_{A,\epsilon}}F)$ is an $\mathsf{RCD}(0,N_{0}+1+D)$-space, where $D=\dim(G)$, and $G_{0}$ acts by measure preserving isometries via the diagonal action $\Delta\colon G\times([0,\infty)\times_{f_{A,\epsilon}}F)$ given by
@@
Thus by [25, Theorem 1.1] the quotient space $(G\times([0,\infty)\times_{f_{A,\epsilon}}F))/\Delta$ is an $\mathsf{RCD}(0,N_{0}+1+D)$ space. But by ?THM? LABEL:MC:_Homeomorphism_rigidity this space is homeomorphic to $X$. ∎
We now consider $N$-warped products of the form $0,2/A]\times_{f_{A}}^{N}F$:
@@
Let $(F,d_{F},\mathfrak{m}_{F})$ be a weighted Finsler manifold which satisfies the $\mathsf{CD}((N-1)K_{F},N)$-condition for $K_{F}>0$ such that, if $N=1$ then $\operatorname{diam}(F)\leq\pi/\sqrt{K_{F}}$. Assume that $K_{F}\geq A^{2}$. Then the $N$-warped product $0,2/A]\times_{f_{A}}^{N}F$ is a $\mathsf{CD}^{\ast}(0,N+1)$-space, and an $\mathsf{RCD}(0,N+1)$-space if $(F,d_{F},\mathfrak{m}_{F})$ satisfies the $\mathsf{RCD}((N-1)K_{F},N)$-condition.
@@
By the previous theorem, the $N$-warped products $0,\infty)\times_{f_{A,\epsilon}}^{N}F$ are $\mathsf{CD}^{\ast}(0,N+1)$-spaces (respectively $\mathsf{RCD}(0,N+1)$-spaces if $F$ is infinitesimally Hilbertian).
We now consider the open subset $0,2/A)\times F/(0,p)\sim(0,q)$. This is the open ball of radius $2/A$ centered at the vertex $\ast$ of the cones, with respect to the metrics $d_{A}$ or $d_{\epsilon}$. Moreover as we showed, it is totally geodesic, and by construction we have that $\mathfrak{m}_{\epsilon}(\partial[0,2/A)\times F/(0,p)\sim(0,q))=0=\mathfrak{m}_{A}(\partial[0,2/A)\times F/(0,p)\sim(0,q))$. Thus we conclude that $0,2/A]\times_{f_{A,\epsilon}}F$ are $\mathsf{CD}^{\ast}(0,N+1)$-spaces (respectively, $\mathsf{RCD}(0,N+1)$-spaces when $F$ is infinitesimally Hilbertian).
Since the $N$-warped product $0,2/A]\times_{f_{A}}^{N}F$ is the pointed measured Gromov-Hausdorff limit of the spaces $0,2/A]\times_{f_{A,\epsilon}}^{N}F$ as $\epsilon\to 0$, and the $\mathsf{CD}^{\ast}$, $\mathsf{RCD}$-conditions are stable under this type of convergence, we obtain the desired result. ∎
@@
Proposition LABEL:prop.productsrescallings (3).
Let $(X,d,\mathfrak{m})$ be an $\mathsf{RCD}(K,N)$. Then for $\alpha,\beta>0$, the space $(X,\alpha d,\beta\mathfrak{m})$ is an $\mathsf{RCD}(\alpha^{-2}K,N)$-space.
We consider an $n$-dimensional closed Riemannian manifold $F$ such that for $0<\kappa_{F}$ we have $\kappa_{F}\leq\mathrm{Ric}(F)$, where $\mathrm{Ric}(F)=\min\{\mathrm{Ric}(x)\mid x\in TM\>\|x\|=1\}$. Thus we have that $(F,d_{F},\mathfrak{m}_{F})$ is a $\mathsf{CD}^{\ast}(\kappa_{F},n)$-space. By taking $\lambda=\sqrt{\kappa_{F}/4(n-1)}$ we have that $(F,\lambda d_{F},\mathfrak{m}_{F})$ is an $\mathsf{CD}^{\ast}((n-1)4,n)$-space. Given that $F$ is a Riemannian manifold, we obtain that $(F,\lambda d_{F},\mathfrak{m}_{F})$ is an $\mathsf{RCD}(0,n)$-space, and also $(F,\mathfrak{m}_{F})$ is a weighted Finsler manifold. Thus we have that $|f_{2,\epsilon}^{\prime}(0)|\leq 2$, and by ?THM? LABEL:T:_cone_over_f_A_sigma_are_CD that $0,\infty)\times_{f_{2,\epsilon}}(F,\lambda d_{F},\mathfrak{m}_{F})$ is an $\mathsf{RCD}(0,n+1)$-space. Since the balls of radius $1$ are geodesic, then by taking the limit as $\epsilon\to 0$ we obtain that that $0,1]\times_{f_{2}}(F,\lambda d_{F},\mathfrak{m}_{F})$ is an $\mathsf{RCD}(0,n+1)$-space.
@@
We now consider gluings along the boundary of (certain quotients of) the $N$-warped products we have considered so far using the functions $f_{A}\colon[0,2/A]\to[0,\infty)$. Consider the cone-like warped product $C=[0,2/A]\times F/(0,p)\sim(0,q)$ with the distance $d_{C}=d_{A}$ and measure $\mathfrak{m}_{C}=f^{n-1}_{A}d\mathfrak{m}_{F}\otimes d\mathcal{L}$, where $(F,d_{F},\mathfrak{m}_{F})$ is an (infinitesimally Hilbertian) weighted Finsler manifold satisfying the $\mathsf{CD}^{\ast}((N-1)K_{F},N)$-condition for some $K_{F}>0$. We consider $G$ equipped with a bi-invariant metric $d_{G}$ and the Haar measure $\mathfrak{m}_{G}$. We take $G\times C$ equipped with the product distance $d_{G}\oplus d_{C}$ and measure $\mathfrak{m}_{G}\otimes\mathfrak{m}_{C}$. Thus, taking $D:=\dim(G)$, by ?THM? LABEL:T:_Special_type_of_cone the space $G\times C$ is an $\mathsf{RCD}(0,N+1+D)$-space.
@@
Consider $G$ an $D$-dimensional compact Lie group, $K\subset G$ a Lie subgroup, and $(F,d_{F},\mathfrak{m}_{F})$ a Riemannian manifold, which satisfies the $\mathsf{RCD}((N-1)K_{F},N)$-condition on which $K$ acts transitively by measure preserving isometries for some $K_{F}>0$. Assume that $\operatorname{diam}(F)\leq\pi/\sqrt{K_{F}}$, if $N=1$. Consider $A>0$ such that $K_{F}>A^{2}$, and the function $f_{A}(x):=\min\{1,Ax\}$. Consider the warped product $C:=[0,2/A]\times_{f_{A}}^{N}F$, and denote by $\Delta$ the diagonal action of $K$ on $G\times C$. Then the space $G\times C/\Delta$ is an $\mathsf{RCD}(0,N+1+D)$-space.
Moreover when we consider a homogeneous space $X=G/H$, equipped with $G$-invariant distance $d$ and measure $\mathfrak{m}$ making it an $\mathsf{RCD}$-space, the measure $\mathfrak{m}$ has to be (a multiple of) the Hausdorff measure as we now show (this has also been independently proved by [47, Theorem 2]).
@@
By [73, Proposition 5.14] we have that $X$ must be isometric to a Riemannian manifold. So, in particular we have that all the tangent spaces must be isometric to $\mathbb{R}^{n}$ for some natural number $n$. Then, by [13, Theorem 1.12] we have then that $\mathfrak{m}=\rho\mathcal{H}^{n}$, where $\mathcal{H}^{n}$ denotes the $n$-dimensional Hausdorff measure. We claim now that for all $g\in G$. $\rho\circ g=\rho$. Let $A\in\mathcal{B}(X)$, $g\in G$, and recall that $g_{\#}\mathcal{H}^{n}=\mathcal{H}^{n}$. As the action is by measure preserving isometries $\mathfrak{m}(A)=g_{\#}\mathfrak{m}(A)$. This implies
$\int_{A}\rho d\mathcal{H}^{n}=\int_{g^{-1}A}\rho d\mathcal{H}^{n}=\int_{A}\rho\circ gd\mathcal{H}^{n}.$
@@
Consider $G_{\pm}$ compact Lie groups, $(P_{\pm},d_{\pm}^{P},\mathfrak{m}_{\pm}^{P})$ be $N_{\pm}^{P}$-dimensional Riemannian manifolds with $\operatorname{Ric}(P_{\pm})\geq K^{P}_{\pm}$, and $(F_{\pm},d_{\pm}^{F},\mathfrak{m}_{\pm}^{F})$ be weighted Riemannian manifolds which are $\mathsf{RCD}(K_{\pm}^{F},N_{\pm}^{F})$ for $N_{\pm}\geq 1$, with $K_{\pm}^{F}>0$ if $N_{\pm}^{F}>1$, and have bounded diameter in the case of $N_{\pm}=1$. Assume that $G_{\pm}$ acts effectively on $P_{\pm}$ and $F_{\pm}$ by measure preserving isometries. Furthermore assume that the balanced product $P_{-}\times_{G_{-}}F_{-}$ is homeomorphic to $P_{+}\times_{G_{+}}F_{+}$. Then the glued space
@@
In the case that $N_{\pm}^{F}>1$, we choose $K_{\pm}>0$ so that $(F_{\pm},d^{F}_{\pm},\mathfrak{m}_{\pm}^{F})$ are $\mathsf{RCD}((N_{\pm}^{F}-1)K_{\pm},N^{F}_{\pm})$, for example by taking $K_{\pm}\leq K_{\pm}^{F}/(N^{F}_{\pm}-1)$. In the case of $N_{\pm}^{F}=1$, we always have that $(F_{\pm},d^{F}_{\pm},\mathfrak{m}_{\pm}^{F})$ are $\mathsf{RCD}((N_{\pm}^{F}-1)K_{\pm},N_{\pm}^{F})=\mathsf{RCD}(0,N_{\pm}^{F})$ for $K_{\pm}>0$. Observe that in the case when either $N^{F}_{+}=1$ or $N^{F}_{-}=1$, then by taking $0<\lambda_{\pm}\leq(\pi/\sqrt{K_{\pm}})/\operatorname{diam}(F_{\pm})$, we have that by rescaling $F_{\pm}$ by $\lambda_{\pm}$, the rescaled space $(F_{\pm},\lambda_{\pm}d^{F}_{\pm},\mathfrak{m}^{F}_{\pm})$ is an $\mathsf{RCD}((N^{F}_{\pm}-1)K_{\pm}/\lambda_{\pm},N^{F}_{\pm})$-space which satisfies $\operatorname{diam}(F_{\pm},\lambda_{\pm}d^{F}_{\pm})\leq\pi/\sqrt{K_{\pm}}$. Alternatively, we can choose $K_{\pm}>0$ such that $\operatorname{diam}(F_{\pm})\leq\pi/\sqrt{K_{\pm}}$. Thus without loss of generality, in case either $N^{F}_{+}=1$ or $N^{F}_{-}=1$, we may assume that $\operatorname{diam}(F_{\pm})\leq\pi/\sqrt{K_{\pm}}$.
We now consider $A_{\pm}>0$ such that $K_{\pm}>A_{\pm}^{2}$. We now consider the following two $\mathsf{RCD}(\bar{K}_{\pm},N^{F}_{\pm}+1+N^{P}_{\pm})$-spaces:
@@
where $\bar{K}_{\pm}=\min\{K_{\pm}^{P},0\}$, and $(d_{A_{\pm}},\mathfrak{m}_{A_{\pm}})$ is the distance and measure of the warped product $0,2/A_{\pm}]\times_{f_{A_{\pm}}}^{N^{F}_{\pm}}F_{\pm}$. The warping function $f_{\pm}$ is given by $f_{\pm}(t)=\min\{1,A_{\pm}t\}$. We denote by $d_{\pm}$ the metric $(d^{P}_{\pm}\otimes d_{A_{\pm}})^{\ast}$, and by $\mathfrak{m}_{\pm}$ the measure $(\mathfrak{m}^{P}_{\pm}\otimes d\mathfrak{m}_{A_{\pm}})_{\ast}$.